# Self-supervised pretraining of vision transformers for animal behavioral analysis and neural encoding

### v1

@@ D.1 Models
The pose estimation models consist of a backbone and a head. The backbone is either a ResNet-50 [16] or a ViT-B/16 [12], both producing feature maps of shape $[N,H,W]$ for a given image, where $N$ denotes the feature dimension and $H$, $W$ denote the height and width of the feature maps. All models employ an identical linear upsampling head that begins with a PixelShuffle layer, reshaping the feature maps to $[N/4,2H,2W]$. These reshaped features then pass through two consecutive 2D convolutional transpose layers with kernel size $(3,3)$ and stride $(2,2)$, doubling the spatial resolution after each layer. The head architecture omits batch normalization and nonlinearities between these layers. The output passes through a 2D softmax function, generating a normalized heatmap for each keypoint.
@@ D.2 Training
We divided the labeled data into training (95%) and validation (5%) sets, with test frames coming from entirely held-out videos. We used a batch size of eight frames. Data augmentations include random crops, rotations, motion blur and histogram equalization. Models were trained for 300 epochs, with validation loss recorded every five epochs. For final evaluation, we selected the model with the lowest validation loss. Training utilized an Adam optimizer [24] with an initial learning rate of 0.001, which was halved at epochs 150, 200, and 250. To facilitate feature learning, we kept the backbone frozen during the first 20 epochs of training before allowing for full end-to-end optimization. The loss function is the mean square error between each predicted heatmap and a ground truth heatmap constructed from labeled data.
@@ E.1 Models
Temporal convolution network
The nonlinear action segmentation model is a dilated TCN [32] with 2 dilation blocks. Each dilation block consists of a sequence of 2 sub-blocks (1D convolution layer → leaky ReLU nonlinearity → dropout with probability=0.10), as well as a residual connection between the input and output of the dilation block. The dilation of the convolutional filters starts with 1 for the first dilation block, then increases by a factor of 2 for each additional dilation block. This results in a larger temporal receptive field as the model gets deeper, allowing for learning of longer range dependencies [78].
@@ E.2 Training
Each video is split into sequences of 500 (IBL) or 100 (CalMS21) frames. We divide the data into training (90%) and validation (10%) sets, with test frames coming from entirely held-out videos. We use a batch size of 16 sequences. Models were trained for 500 epochs using the Adam optimizer [24].
@@ E.4 Temporal convolution network with multi-head attention pooling
We aggregate the patch embeddings from the Beast encoder using a multi-head attention pooling layer [33], which produces a single pooled embedding per frame as input to the TCN. This layer uses a learnable query $S\in\mathbb{R}^{1\times D}$ with patch embeddings $z\in\mathbb{R}^{N\times D}$ as keys and values. To capture motion-related features, we further concatenate the frame-to-frame difference of the pooled embeddings as additional input to the TCN (Fig. 8). We fixed the number of attention heads in the pooling layer to 8 for all models.
@@
The pooling layer and TCN classifier were trained jointly for 200 epochs on CalMS21 and 100 epochs on IBL using the Adam optimizer [24] with an initial learning rate of $1\mathrm{e}{-3}$. The epoch counts were determined through a separate experiment that withheld a subset of validation videos and monitored the validation F1 score until convergence. Training followed a cosine-annealing schedule with warm restarts [36] configured with $T_{0}=34$, $T_{mult}=2$ and $\eta_{min}=5\mathrm{e}{-5}$. We used 6 or 8 NVIDIA A40 GPUs for the training, each with a batch size of 2, giving an effective batch size of 12 or 16. The sequence length was fixed at 500. Due to the increased compute required to train these models, we fixed the TCN hyperparameters to be those found for the CLS-based model for each dataset.
@@
Table 13: Action segmentation results on IBL and CalMS21 datasets across feature types and models. All ViT-based models use frozen a frozen backbone; “CLS” indicates models trained on the global CLS embeddings, while “patch” indicates models trained with a multi-head attention pooling layer applied to the patch embeddings. “IN” refers to a model pretrained with ImageNet weights, “IN+PT” refers to models that are initialized with ImageNet-pretrained weights then further pretrained on experiment-specific data.
@@
Dataset Features Linear TCN Features Features, $\Delta$ Features Features Features, $\Delta$ Features IBL Keypoints $0.54\pm 1.4\mathrm{e}{-3}$ $0.55\pm 1.5\mathrm{e}{-3}$ $0.86\pm 1.4\mathrm{e}{-3}$ $0.88\pm 2.2\mathrm{e}{-3}$ PCA $0.54\pm 4.6\mathrm{e}{-3}$ $0.55\pm 6.3\mathrm{e}{-3}$ $0.64\pm 1.0\mathrm{e}{-2}$ $0.71\pm 2.8\mathrm{e}{-3}$ VIT-M (IN) (CLS) $0.68\pm 1.4\mathrm{e}{-3}$ $0.68\pm 7.0\mathrm{e}{-4}$ $0.78\pm 7.4\mathrm{e}{-3}$ $0.79\pm 2.7\mathrm{e}{-3}$ VIT-M (IN+PT) (CLS) $0.74\pm 2.5\mathrm{e}{-3}$ $0.72\pm 2.0\mathrm{e}{-3}$ $0.78\pm 4.0\mathrm{e}{-3}$ $0.80\pm 4.6\mathrm{e}{-3}$ Beast (IN+PT) (CLS) $0.70\pm 7.9\mathrm{e}{-3}$ $0.69\pm 2.5\mathrm{e}{-3}$ $0.80\pm 3.0\mathrm{e}{-3}$ $0.81\pm 6.9\mathrm{e}{-4}$ VIT-M (IN) (patch) - - - $0.84\pm 3.7\mathrm{e}{-3}$ VIT-M (IN+PT) (patch) - - - $0.85\pm 3.6\mathrm{e}{-3}$ Beast (IN+PT) (patch) - - - $0.87\pm 5.1\mathrm{e}{-3}$ CalMS21 SimBA [14] $0.75\pm 5.1\mathrm{e}{-4}$ $0.53\pm 8.1\mathrm{e}{-3}$ $0.78\pm 3.8\mathrm{e}{-3}$ $0.79\pm 2.9\mathrm{e}{-3}$ TREBA [60] $0.29\pm 1.4\mathrm{e}{-3}$ $0.30\pm 1.5\mathrm{e}{-3}$ $0.70\pm 4.6\mathrm{e}{-3}$ $0.72\pm 7.4\mathrm{e}{-3}$ PCA $0.10\pm 3.1\mathrm{e}{-3}$ $0.10\pm 3.2\mathrm{e}{-3}$ $0.16\pm 5.1\mathrm{e}{-3}$ $0.18\pm 4.5\mathrm{e}{-3}$ VIT-M (IN) (CLS) $0.50\pm 2.2\mathrm{e}{-2}$ $0.52\pm 1.3\mathrm{e}{-3}$ $0.53\pm 1.2\mathrm{e}{-2}$ $0.60\pm 2.8\mathrm{e}{-3}$ VIT-M (IN+PT) (CLS) $0.60\pm 5.5\mathrm{e}{-3}$ $0.60\pm 1.1\mathrm{e}{-3}$ $0.60\pm 9.1\mathrm{e}{-3}$ $0.65\pm 2.2\mathrm{e}{-3}$ Beast (IN+PT) (CLS) $0.53\pm 4.6\mathrm{e}{-3}$ $0.51\pm 4.0\mathrm{e}{-2}$ $0.58\pm 5.2\mathrm{e}{-3}$ $0.63\pm 2.7\mathrm{e}{-3}$ VIT-M (IN) (patch) - - - $0.74\pm 2.9\mathrm{e}{-3}$ VIT-M (IN+PT) (patch) - - - $0.82\pm 9.5\mathrm{e}{-3}$ Beast (IN+PT) (patch) - - - $0.81\pm 7.7\mathrm{e}{-3}$
Appendix F Broader Impacts

### v2

@@ D.1 Models
The pose estimation models consist of a backbone and a head. The backbone is either a ResNet-50 [He et al., 2016] or a ViT-B/16 [Dosovitskiy et al., 2020], both producing feature maps of shape $[N,H,W]$ for a given image, where $N$ denotes the feature dimension and $H$, $W$ denote the height and width of the feature maps. All models employ an identical linear upsampling head that begins with a PixelShuffle layer, reshaping the feature maps to $[N/4,2H,2W]$. These reshaped features then pass through two consecutive 2D convolutional transpose layers with kernel size $(3,3)$ and stride $(2,2)$, doubling the spatial resolution after each layer. The head architecture omits batch normalization and nonlinearities between these layers. The output passes through a 2D softmax function, generating a normalized heatmap for each keypoint.
@@ D.2 Training
We divided the labeled data into training (95%) and validation (5%) sets, with test frames coming from entirely held-out videos. We used a batch size of eight frames. Data augmentations include random crops, rotations, motion blur and histogram equalization. Models were trained for 300 epochs, with validation loss recorded every five epochs. For final evaluation, we selected the model with the lowest validation loss. Training utilized an Adam optimizer [Kingma and Ba, 2014] with an initial learning rate of 0.001 for the ResNet and $5\mathrm{e}{-5}$ for the transformers, which was halved at epochs 150, 200, and 250. To facilitate feature learning, we kept the backbone frozen during the first 20 epochs of training before allowing for full end-to-end optimization. The loss function is the mean square error between each predicted heatmap and a ground truth heatmap constructed from labeled data.
D.3 Pretrained backbones
We test several additional ViT backbones on pose estimation to validate our Beast results:
These backbones are trained using the same procedure as the ResNet-50 and Beast models (see above). We find other pretrained backbones mostly outperform the ResNet-50 baseline (Fig. 9). SAM is generally the least performant backbone. DINOv2 consistently outperforms DINO across all datasets, and Beast achieves the lowest pixel error in most cases (only outperformed by DINOv2 in the Mirror-fish dataset). These results demonstrate Beast’s experiment-specific pretraining framework can surpass state-of-the-art general purpose vision foundation models for pose estimation.
Figure 9: Pose estimation of fine-tuned vision foundation model backbones. We evaluated ResNet-50 (pretrained on AP10K), SAM, DINO, DINOv2, and Beast backbones on pose estimation datasets. We evaluate the these models using pixel error at various ensemble standard deviation thresholds, with values in the table representing the percentage of keypoints at the chosen threshold. Smaller values indicate smaller but more challenging subsets of keypoints (see main text).
@@ D.3 Pretrained backbones
D.4 DeepLabCut baseline
For the DeepLabCut baseline (version 3.0.0) we trained models using an ImageNet-pretrained ResNet-50 backbone. To properly isolate differences between DeepLabCut and Lightning Pose algorithms, we matched training frames, batch size, learning rate schedule, and number of epochs. For all other hyperparameters we used the DeepLabCut package defaults (e.g., data augmentation). We train models using three different train/val data splits, and ensure these splits exactly match those used for the Lightning Pose models.
Lightning Pose outperforms DeepLabCut on the IBL, Mirror-mouse, and Mirror-fish datasets, while DeepLabCut outperforms Lightning Pose on CRIM13 (Fig. 11). Notably, our pretrained Beast backbones outperform both DeepLabCut and Lightning Pose across all four datasets.
Figure 11: Pose estimation of DeepLabCut (DLC). Figure conventions as in Fig. 9
@@ E.1 Models
Temporal convolution network.
The nonlinear action segmentation model is a dilated TCN [Lea et al., 2016] with 2 dilation blocks. Each dilation block consists of a sequence of 2 sub-blocks (1D convolution layer → leaky ReLU nonlinearity → dropout with probability=0.10), as well as a residual connection between the input and output of the dilation block. The dilation of the convolutional filters starts with 1 for the first dilation block, then increases by a factor of 2 for each additional dilation block. This results in a larger temporal receptive field as the model gets deeper, allowing for learning of longer range dependencies [Yu and Koltun, 2015].
@@ E.2 Training
Each video is split into sequences of 500 (IBL) or 100 (CalMS21) frames. We divide the data into training (90%) and validation (10%) sets, with test frames coming from entirely held-out videos. We use a batch size of 16 sequences. Models were trained for 500 epochs using the Adam optimizer [Kingma and Ba, 2014].
@@ E.4 Temporal convolution network with multi-head attention pooling
We aggregate the patch embeddings from the Beast encoder using a multi-head attention pooling layer [Lee et al., 2019], which produces a single pooled embedding per frame as input to the TCN. This layer uses a learnable query $S\in\mathbb{R}^{1\times D}$ with patch embeddings $z\in\mathbb{R}^{N\times D}$ as keys and values. To capture motion-related features, we further concatenate the frame-to-frame difference of the pooled embeddings as additional input to the TCN (Fig. 12). We fixed the number of attention heads in the pooling layer to 8 for all models.
@@
The pooling layer and TCN classifier were trained jointly for 200 epochs on CalMS21 and 100 epochs on IBL using the Adam optimizer [Kingma and Ba, 2014] with an initial learning rate of $1\mathrm{e}{-3}$. The epoch counts were determined through a separate experiment that withheld a subset of validation videos and monitored the validation F1 score until convergence. Training followed a cosine-annealing schedule with warm restarts [Loshchilov and Hutter, 2016] configured with $T_{0}=34$, $T_{mult}=2$ and $\eta_{min}=5\mathrm{e}{-5}$. We used 6 or 8 NVIDIA A40 GPUs for the training, each with a batch size of 2, giving an effective batch size of 12 or 16. The sequence length was fixed at 500. Due to the increased compute required to train these models, we fixed the TCN hyperparameters to be those found for the CLS-based model for each dataset.
@@
Table 15: Action segmentation results on IBL and CalMS21 datasets across feature types and models. All ViT-based models use frozen a frozen backbone; “CLS” indicates models trained on the global CLS embeddings, while “patch” indicates models trained with a multi-head attention pooling layer applied to the patch embeddings. “IN” refers to a model pretrained with ImageNet weights, “IN+PT” refers to models that are initialized with ImageNet-pretrained weights then further pretrained on experiment-specific data. We report the mean and standard deviation of F1 on test data across five random train/val splits.
@@
Dataset Features Linear TCN Features Features, $\Delta$ Features Features Features, $\Delta$ Features IBL Keypoints $0.54\pm 1.4\mathrm{e}{-3}$ $0.55\pm 1.5\mathrm{e}{-3}$ $0.86\pm 1.4\mathrm{e}{-3}$ $\mathbf{0.88\pm 2.2\mathrm{e}{-3}}$ PCA $0.54\pm 4.6\mathrm{e}{-3}$ $0.55\pm 6.3\mathrm{e}{-3}$ $0.64\pm 1.0\mathrm{e}{-2}$ $0.71\pm 2.8\mathrm{e}{-3}$ VIT-M (IN) (CLS) $0.68\pm 1.4\mathrm{e}{-3}$ $0.68\pm 7.0\mathrm{e}{-4}$ $0.78\pm 7.4\mathrm{e}{-3}$ $0.79\pm 2.7\mathrm{e}{-3}$ VIT-M (IN+PT) (CLS) $\mathbf{0.74\pm 2.5\mathrm{e}{-3}}$ $\mathbf{0.72\pm 2.0\mathrm{e}{-3}}$ $0.78\pm 4.0\mathrm{e}{-3}$ $0.80\pm 4.6\mathrm{e}{-3}$ Beast (IN+PT) (CLS) $0.70\pm 7.9\mathrm{e}{-3}$ $0.69\pm 2.5\mathrm{e}{-3}$ $\mathbf{0.80\pm 3.0\mathrm{e}{-3}}$ $0.81\pm 6.9\mathrm{e}{-4}$ DINOv2 (patch) - - - $0.77\pm 2.7\mathrm{e}{-3}$ VIT-M (IN) (patch) - - - $0.84\pm 3.7\mathrm{e}{-3}$ VIT-M (IN+PT) (patch) - - - $0.85\pm 3.6\mathrm{e}{-3}$ ViT-C (IN+PT) (patch) - - - $0.79\pm 4.6\mathrm{e}{-3}$ Beast (IN+PT) (patch) - - - $0.87\pm 5.1\mathrm{e}{-3}$ CalMS21 SimBA [Goodwin et al., 2024] $\mathbf{0.75\pm 5.1\mathrm{e}{-4}}$ $0.53\pm 8.1\mathrm{e}{-3}$ $\mathbf{0.78\pm 3.8\mathrm{e}{-3}}$ $0.79\pm 2.9\mathrm{e}{-3}$ TREBA [Sun et al., 2021b] $0.29\pm 1.4\mathrm{e}{-3}$ $0.30\pm 1.5\mathrm{e}{-3}$ $0.70\pm 4.6\mathrm{e}{-3}$ $0.72\pm 7.4\mathrm{e}{-3}$ PCA $0.10\pm 3.1\mathrm{e}{-3}$ $0.10\pm 3.2\mathrm{e}{-3}$ $0.16\pm 5.1\mathrm{e}{-3}$ $0.18\pm 4.5\mathrm{e}{-3}$ VIT-M (IN) (CLS) $0.50\pm 2.2\mathrm{e}{-2}$ $0.52\pm 1.3\mathrm{e}{-3}$ $0.53\pm 1.2\mathrm{e}{-2}$ $0.60\pm 2.8\mathrm{e}{-3}$ VIT-M (IN+PT) (CLS) $0.60\pm 5.5\mathrm{e}{-3}$ $\mathbf{0.60\pm 1.1\mathrm{e}{-3}}$ $0.60\pm 9.1\mathrm{e}{-3}$ $0.65\pm 2.2\mathrm{e}{-3}$ Beast (IN+PT) (CLS) $0.53\pm 4.6\mathrm{e}{-3}$ $0.51\pm 4.0\mathrm{e}{-2}$ $0.58\pm 5.2\mathrm{e}{-3}$ $0.63\pm 2.7\mathrm{e}{-3}$ DINOv2 (patch) - - - $0.68\pm 4.4\mathrm{e}{-3}$ VIT-M (IN) (patch) - - - $0.74\pm 2.9\mathrm{e}{-3}$ VIT-M (IN+PT) (patch) - - - $\mathbf{0.82\pm 9.5\mathrm{e}{-3}}$ ViT-C (IN+PT) (patch) - - - $0.53\pm 2.9\mathrm{e}{-2}$ Beast (IN+PT) (patch) - - - $0.81\pm 7.7\mathrm{e}{-3}$
Appendix F Self-supervised learning methods for animal behavior
Self-supervised learning (SSL) techniques from computer vision have increasingly been adapted to the study of animal behavior. These approaches use SSL to extract useful feature representations from pose, image, or video data, which are then applied to downstream tasks such as action segmentation.
Pose-based approaches.
Trajectory Embedding for Behavior Analysis (TREBA) [Sun et al., 2021b] employs a multi-task self-supervised framework that uses trajectory reconstruction as its primary objective through Trajectory Variational Autoencoders [Co-Reyes et al., 2018, Zhan et al., 2020]. TREBA additionally requires the TVAE embedding to decode various auxiliary tasks consisting of simple data transformations designed by domain experts, with the resulting embeddings serving as input to downstream action segmentation models. Variational Animal Motion Embedding (VAME) [Luxem et al., 2022] uses a sequential variational autoencoder to embed pose sequences into a latent space by reconstructing both current and subsequent time steps. Unlike TREBA, VAME applies clustering to the learned embeddings, creating a fully unsupervised action segmentation pipeline. ContrastivePose [Zhou et al., 2022] leverages geometric augmentations (flipping, rotation, translation) of pose coordinates to generate positive pairs for contrastive learning, followed by fine-tuning on action segmentation tasks. Bootstrap Across Multiple Scales (BAMS) [Azabou et al., 2023] employs dual temporal convolutional networks with different receptive field sizes to create complementary short- and long-term embedding spaces. BAMS introduces a novel training objective requiring prediction of future action distributions rather than specific action sequences, with validation on the MABe benchmark [Sun et al., 2023] across multiple tasks including action segmentation and mouse strain classification.
Image and video-based approaches.
Selfee [Jia et al., 2022] constructs composite RGB frames from 3-frame grayscale video sequences (assigning each frame to a separate color channel) and applies standard image-based contrastive learning techniques, demonstrating effectiveness on action segmentation and anomaly detection. Mueller et al. [2025] adapt a pretrained V-JEPA model [Bardes et al., 2023], an SSL approach specialized for video understanding, through domain-adaptive pretraining on primate behavior datasets, validating their approach on behavior recognition tasks. Similarly, Animal-JEPA [Zheng et al., 2024] modifies the V-JEPA training strategy with domain-specific masking techniques and validates on mouse behavior classification tasks.
Distinction from prior work.
While these methods share conceptual similarities with Beast, our approach is distinguished by its general frame-based training objectives and comprehensive evaluation across neural activity prediction and pose estimation tasks, in addition to the standard action segmentation task. Furthermore, since the pose-based methods described above rely on pose estimates as input, Beast could potentially enhance their performance by providing higher-quality pose estimation as a preprocessing step.
Appendix G Beast workflow
We give an overview of the Beast workflow for a new user, highlighting steps where Beast enhances a traditional workflow.
Stage 1: Data collection
  • Collect behavioral videos
  • Optionally collect simultaneous neural recordings
The use of Beast does not affect this step of the workflow.
Stage 2: Beast pretraining
This step will be skipped in a traditional workflow.
  • Extract frames from unlabeled videos (typically $\sim$100K frames)
  • Train Beast transformer backbone on these frames ($\sim$12 hours on 8 GPUs)
  • This step does not require manual annotation
Stage 3: Downstream applications
Option A: Pose estimation
  • Annotate 100-1000 frames from 5-50 videos with keypoints (can be different from pretraining set)
  • Fine-tune pose estimation model using Beast backbone
    Key advantage: improved performance over existing backbones (Fig. 3)
  • Run inference on new videos to extract keypoints
Option B: Action segmentation
  • Annotate 1000-5000 frames per behavior, ideally across 5-10 videos, with action labels (independent of pose annotation)
  • Fine-tune action segmentation model using Beast backbone
    Key advantage 1: skip pose estimation pipeline entirely
    Key advantage 2: equivalent or better performance compared to pose estimates (Fig. 4)
  • Run inference on new videos to extract frame-by-frame actions
Option C: Neural encoding
  • No manual annotation (neural activity provides the “labels”)
  • Fine-tune neural encoding model using Beast backbone (video input, neural activity output)
    Key advantage 1: skip pose estimation pipeline entirely
    Key advantage 2: improved performance over existing behavioral features (Fig. 2)
  • Predict neural activity from video
Appendix H Broader Impacts